If 3 times the smaller of two consecutive integers is added to four tmes the larger the result is 130. Find the numbers.
Answer by ikdeep(189)

Let the first or smaller consecutive integer = x
Let the second or larger consecutive integer = x+1
Now according to equation,,,we get
3(x) + 4(x+1) = 130
3x + 4x + 4 = 130
7x + 4 = 130
subtracting 4 from both sides we get ...
7x = 126
divide both sides by 7 and we get
x = 18
so smaller consecutive integer = x = 18
and Larger consecutive integer = x + 1 = 19
hope this will help you
Please feel free to revert back for any further queries.